If DV/Ds * v= a  and v=0 then what a is zero or non zero

Grade:11

1 Answers

Arun
4 years ago
if a = v.(dv/dx) ,
put v = dx/dt ,
=> a = (dx/dt)*(dv/dx)
=> a=dv/dt ;

the answer to the question u seek lies in the fundamentals of calculus. The formula a= v*dv/dx is actually derived from the definition of a=dv/dt.
v = dx/dt
v • dv/dx = dx/dt • dv/dx = dv/dt = a
here, when we are cancelling out the "dx" term, we are assuming that it is non-zero. However, in the throwing ball problem, "dx" at the top most point is zero. Hence you are getting erroneous answer.